OpenStudy (anonymous):

what are the foci of the ellipse given by the equation 100x^2+64y^2=6,400

OpenStudy (jdoe0001):

hmmm do you know an ellipse's equation?

OpenStudy (anonymous):

could you remind me?

OpenStudy (jdoe0001):

\(\bf 100x^2+64y^2=6,400\implies \cfrac{100x^2}{6400}+\cfrac{64y^2}{6400}=1\implies \cfrac{x^2}{64}+\cfrac{y^2}{100}=1\\ \quad \\ \cfrac{x^2}{8^2}+\cfrac{y^2}{10^2}=1\implies \cfrac{(x-0)^2}{8^2}+\cfrac{(y-0)^2}{10^2}=1\)

OpenStudy (jdoe0001):

the bigger denominator is under the "y", so that means the ellipse is moving vertically so the foci will be \(\bf "c"\) distance from the center of it, over the y-axis \(\bf c=\sqrt{a^2-b^2}\)

OpenStudy (jdoe0001):

"a" component is always the BIGGER denominator in an ellipse "b" is the smaller one

OpenStudy (jdoe0001):

anyhow, to avoid any ambiguities from the squared form a = 10 b = 8

OpenStudy (anonymous):

\[0,\pm36\]

OpenStudy (anonymous):

is that correct i followed you formula

OpenStudy (jdoe0001):

close, \(\bf c = \sqrt{a^2-b^2}\implies c=\sqrt{100-64}\implies c = \sqrt{36}\)

OpenStudy (anonymous):

so the \[\pm36\] comes first?

OpenStudy (jdoe0001):

nope, you're correct, you just forgot to take square root of 36

OpenStudy (jdoe0001):

\(\bf c = \sqrt{a^2-b^2}\implies c=\sqrt{100-64}\implies c = \sqrt{36}\\ \quad \\ c = 6\qquad \qquad (0,\pm 6)\)

OpenStudy (anonymous):

ohhhhhh gotchaaaaaa